OpenStudy (anonymous):

write each function as either a sum/difference of logs. express exponents as factors ln(x/e^3)

OpenStudy (anonymous):

\[\ln(x)-\ln(e ^{3}) \rightarrow \ln(x) - 3\ln(e) \rightarrow \ln(x) -3\]

OpenStudy (anonymous):

thank you very much :) the last one is the answer?